OpenStudy (anonymous):

what is 12 * 3 - x If x = 18/3

Mehek (mehek14):

\(12*3-\dfrac{18}{3}\) what's \(12*3=?\)

Mehek (mehek14):

@MarenaRose

OpenStudy (setsuna-yuregeshi):

What is 12 times 3?? 36, so now, what is: \[\frac{ 36 }{ 1 } - \frac{ 18 }{ 3 }\] How do you subtract those two fractions?